OpenStudy (anonymous):

Can someone help me? My problem says, Multiply (sqrt10 -9)^2

OpenStudy (anonymous):

you gotta work this out: (sqrt10-9)(sqrt10-9)

OpenStudy (anonymous):

Ok. Tryin now.

OpenStudy (anonymous):

I get sqrt 100-81.

OpenStudy (anonymous):

you should get : 10-18sqrt10+81

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

I know the answer, 91-18 sqrt 10. Don't know how to get it.

OpenStudy (anonymous):

use the identity (a-b)^2 to check your answers.... (a-b)^2=a^2+b^2-2ab for this question simply multiply (sq. root 10-9)(sq. root10-9) i.e, 10-9sq. root 10-9 sq.root 10+81 =91-18 sq. root 10

OpenStudy (anonymous):

How?